# Evolutionary Computation in the Era of Large Language Model: Survey and Roadmap

### v1

@@ IV Integrated Synergy and Applications of LLM and EA
In recent years, the synergy between LLM and EA has attracted increasing attention. They combine the strengths of LLM and EA to enhance performance in various downstream applications. For instance, LLM can be used for natural language understanding and provide domain knowledge, while EA serves as an iterative search framework to find better solutions. The integration of LLM and EA can lead to a powerful domain-knowledge-based optimization approach, which is problem-dependent and designed for specific problems. In this section, we will provide a review from a problem-based perspective and discuss the collaborative effects of LLM and EA in several popular downstream applications.
IV-A Neural Architecture Search
Throughout the years, EAs have found widespread application in NAS [64, 65], revolutionizing the way we design and optimize neural networks. Their ability to simulate natural selection and iteratively improve architectures has proven invaluable in the quest for efficient and high-performing models. However, as we delve into the realm of cutting-edge advancements, a new player has emerged onto the NAS scene - LLMs [68, 69, 70]. These powerful models, such as GPT-4 [4], have transcended their traditional language generation capabilities and are now being harnessed to tackle the intricate challenges of architecture search. With their immense computational capacity and deep understanding of language and context, LLMs bring a fresh perspective to NAS. They possess the potential to offer novel insights and innovative solutions, empowering researchers and engineers to explore uncharted territories of network design.
Yu et al. proposed a NAS method called GPT-NAS [71] that utilizes a GPT model [72] to guide the search of neural architectures. In GPT-NAS, neural architectures are encoded as inputs for the GPT model using a defined encoding strategy. Then the GPT model is pre-trained and fine-tuned on neural architecture datasets to introduce prior knowledge into the search process. In the architecture search, GPT-NAS uses an EA with operations like crossover and mutation to optimize individuals and obtain offspring with better performance. The fine-tuned GPT model is then used to effectively predict excellent new architectures based on previous structural information, thereby reconstructing the sampled architectures to guide and optimize the entire search process. In this method, the EA and GPT achieve complementary advantages. The EA is strong in search and optimization. But directly using the EA may be limited by its variation capability and the huge search space, making it difficult to truly mine excellent architectures. By reorganizing the structures on the basis of individuals evolved by GPT, richer prior knowledge can be introduced to optimize the network architectures, effectively reducing the actual search space and guiding the search towards high-performance areas.
Nasir et al. leveraged the code generation ability of large models to help NAS [73]. LLMatic proposed by them does not directly search in the structure representation space, but uses the code of neural network structures as search points. Specifically, LLMatic establishes two archives: a network archive that stores trained network models, and a prompt archive that stores prompts and temperature parameters for generating networks. LLMatic employs Quality-Diversity (QD) algorithms [74] to perform the search, utilizing the repeated process of generation, training and selection to continuously optimize the quality and diversity of networks. The large model CodeGen [75] is used to complete crossover and mutation in the evolutionary process: in the mutation operation, CodeGen makes subtle modifications to the selected network based on the mutation prompt. In the crossover operation, CodeGen partially matches the selected multiple networks based on the crossover prompt. Chen et al. took a similar approach by using the PALM [8] as the crossover and mutation operators for the NAS task in the evolutionary search process. Moreover, they further improved the LLM’s ability of generating candidate architectures through soft prompt tuning during NAS [76]. Empirical studies showed that this method can find neural network architectures with better performance and smaller size. In addition to the above study, some papers [69, 70] do not explicitly mention using an EA framework, but employ LLM as an operator to generate new architectures.
Unlike previous works that used LLM in the NAS search process, Jawahar et al. utilized LLM to build performance predictors [77]. Specifically, they designed predictor prompts (including task description, architecture parameter definitions, examples, etc.) to define the prediction task. The prompts and test architectures would be input into the LLM (such as GPT-4), and the LLM could make predictions based on the architecture knowledge it learned during pre-training by understanding the complex relationships between architectures and performances. Since LLMs learned architecture-related knowledge from a large amount of literature, and this type of predictor is designed simply without needing to train an over-parameterized network, the deployment cost is low. Therefore, the performance predictor based on LLM is an efficient and accurate method.
IV-B Code Generation
@@
LLMs and EAs have both shown promise in automating code generation. LLMs can be trained on vast amounts of publicly available source code to gain a broad understanding of programming concepts and patterns [78, 79]. However, their generation abilities are limited by their training data distribution. EAs, on the other hand, are capable of open-ended search through program spaces. But traditional mutation operators used in genetic programming struggle to propose high-quality changes in a way that mimics how human programmers intentionally modify code. The joining of EAs and LLMs has opened up more opportunities for code generation.
IV-B1 Universal Code Generation
Lehman et al. applied LLMs into genetic programming (GP), called evolution through large models (ELM) [39]. ELM utilizes a small number of initially hand-written programs as seeds, combining the MAP-Elites search algorithm [80, 81] with an LLM-based intelligent mutation operation to generate a large amount of functionally rich programs as design examples. This method improves the search efficiency of GP algorithms since the LLM trained on code promote a more intelligent and effective mutation operator and reducing the probability of random mutations generating meaningless code. Considering the LLM may achieve different performance improvements in different domains, these programs are used as a training set to train code generation LLMs in order to guide the LLM to output better code. Furthermore, by splicing new conditional inputs onto the LLM, it can teach the LLM to generate conditionally. This model will be fine-tuned through reinforcement learning to make the model able to generate appropriate outputs based on observed conditions. Lehman et al. also explored the effectiveness of ELM in the Sodarace domain [82] with GPT-3 [7].
@@ IV-B1 Universal Code Generation
IV-B2 Domain-specific Code Generation
Chen et al. leverage the code generation capabilities of LLMs for customized data cleaning tasks [37], addressing the limitations of traditional inflexible data cleaning tools and the heavy workload of manual development. Their proposed SEED algorithm takes the JSON configuration file describing the task information as input, and utilizes LLMs to generate the data cleaning pipeline, including not only code generation for each module, but also data cleaning plan generation, module partitioning, data annotation, tool selection for each step, as well as prompt generation - all of which are completed by LLMs.
Ma et al. proposed an algorithm called EUREKA for designing reward functions in reinforcement learning [36]. It utilizes the zero-shot code generation and in-context improvement capabilities of LLM to perform evolutionary optimization of reward codes. Specifically, EUREKA first uses the environment source code as context to realize zero-shot generation of reward functions by GPT-4. Through a simplified evolutionary search, it continuously proposes new reward candidate sets and utilizes in-context improvement based on textual reward feedback to continuously optimize the best reward function. At the same time, EUREKA executes multiple independent searches and allows human-provided reward feedback to guide learning behaviors more aligned with human preferences and find better global optimal solutions.
Fu et al. proposed GPT4AIGChip, a framework that leverages LLM and EA to automate the design of AI accelerators, aiming to address the issue that existing tools require extensive hardware expertise [83]. GPT4AIGChip uses natural language instead of hardware-specific languages to automate AI accelerator design. It iterates the following steps: First, the search engine iteratively generates the next design task and its description using evolutionary search algorithms [84]. The prompt generator selects the two most relevant demonstration-code pairs based on the design description, and combines them with the description to form a prompt provided to the GPT-4. The LLM generates code based on the prompt utilizing in-context learning. This code then undergoes validation and optimization using High-Level Synthesis tools, and performance metrics are obtained on FPGA. The results are fed back to the search engine for the next round of search.
Researchers from DeepMind and other institutions introduced a method called FunSearch [30], which leverages the code generation capabilities of LLMs and EAs for mathematical and algorithmic discovery. In the FunSearch method, the GPT-3 is used to generate initial candidate solution programs, while the EA works to optimize and evolve these programs based on scores from a systematic evaluator. The EA varies the programs through operations like inserting, deleting or replacing statements. Over multiple rounds of ”trial-and-error learning”, the FunSearch method is able to surpass existing best known results by discovering new and improved solutions through iteratively evolving the candidate programs. Compared to just outputting answers directly, FunSearch discovers more interpretable solution programs that can enable expert feedback to further push the boundaries of what’s possible.
IV-B3 Security of Generated Code
@@
In addition to directly generating code, some studies have focused on the security of generated code. Wu et al. concerned about how to evaluate whether the generated code is vulnerable to attacks [85]. They proposed DeceptPrompt, which can generate adversarial natural language prefixes/suffixes to drive code generation models to produce functionally correct code containing specified vulnerabilities. DeceptPrompt continuously optimizes the prefixes and suffixes using genetic algorithms [86], with LLM serving as the mutation operator. The fitness function is separately designed for the functionally correct part and the vulnerable part to guide the generated code to both retain functionality and contain the specified vulnerabilities. Empirical studies have shown that DeceptPrompt can successfully attack various popular code generation models. Tao et al. further considered methods to enhance the security of generated code [87]. They proposed a simple idea of combining LLM with Grammar-Guided Genetic Programming (G3P) system [88], which enforces that any synthesized program adheres to BNF syntax, thus promoting the development/repair of incorrect programs and reducing the chances of security threats. Specifically, LLM is primarily employed to generate the initial code population in G3P, where the initial population is mapped to a predefined BNF syntax program before evolutionary search, allowing for improvements of these programs using program synthesis benchmarks. Afterwards, these programs will be handed over to G3P and evolved into better code.
IV-C Software Engineering
@@
Due to the promising performance of LLM and EA in code generation tasks, some studies have further applied them to practical problems in software engineering, including subtakss software optimization, software testing, and software project planning.
IV-C1 Software Optimization
@@
Genetic Improvement (GI) [31] is a technique for automatically software optimization based on the ideas of evolutionary computation simulating the principles of biological evolution. Kang et al. leverage the advantages of LLM in code understanding and generation to improve the shortcomings of blind mutation in traditional GI [89]. The usage is similar to the code generation steps in Section IV-B, firstly letting the LLM improve the code for specific optimization objectives (such as time efficiency and memory consumption), then injecting the mutations generated by LLM into the candidate pool of GI, and continuously performing evolutionary computation operations until termination to obtain optimized code.
IV-C2 Software Testing
@@
EAs have extensive applications in software testing techniques, where they are employed to search for effective test cases within the testing space to uncover errors and defects in software systems [90, 91]. Recently, some studies combine the code generation capability of LLMs and search capacity of EAs, and apply them to software testing. Deng et al. proposed TitanFuzz [29] to test for bugs in deep learning libraries. The core idea of TitanFuzz is to directly leverage LLMs that have been pre-trained on tens of billions of code snippets, which implicitly learned the syntax and API constraints. It utilizes the generative ability of Codex [78] for seed generation and the infilling ability of InCoder [92] for mutation generation, to automatically produce a large number of input programs that meet the requirements for testing. The fitness function considers both the depth of the dataflow graph obtained from static analysis and the number of unique APIs called in the program, as well as the number of repeated API calls (as a penalty). This encourages the generation of programs with more complex API usage and richer API interactions. Empirical evaluation proves that TitanFuzz can better cover libraries and complex API interactions in this way, thus discovering more elusive bugs that are difficult to find. Lemieux et al. took a more direct approach [93]. They first used conventional evolutionary search (e.g., MOSA [94]) and monitored its coverage progress. When the algorithm reached a coverage plateau (state where further mutation of test cases finds it difficult to improve coverage [95, 96]), it would identify functions with low coverage in the module, and use these low-coverage functions as hints to request the LLM (e.g., CodeX [78]) to generate test cases. Afterwards, on the basis of the test cases generated by the LLM, the EA continued exploration. This method can significantly improve the coverage of test samples.
IV-C3 Software Project Planning
@@
Story points are a commonly used method in software project planning and cost estimation to gauge the amount of work required to complete a user story [97]. One way to leverage LLMs for this task is through few-shot learning, by providing some labeled examples of past user stories and their estimated points to improve the model’s ability to estimate new stories. However, the choice and number of examples can impact the estimation effectiveness. Tawosi et al. propose to optimize the example selection through a multi-objective approach [98]. They use NSGA-II [53] to simultaneously optimize three objective functions - the sum of absolute errors, the confidence interval of the error distribution, and the number of examples. This searches for the Pareto front of user story example sets. The genetic algorithm explores the trade-off between estimation accuracy and complexity, as measured by the objective functions. The optimized Pareto front provides decision makers with different accuracy-complexity choice options when selecting an example set for few-shot learning with LLMs to estimate story points.

### v3

@@ IV Applications Driven by
In recent years, the synergy between LLM and EA has attracted increasing attention. Researchers combine the strengths of LLM and EA to enhance performance in various downstream applications, as shown in Table V. This section provides a review from a problem-based perspective and discuss the collaborative effects of LLM and EA in several popular downstream applications.
IV-A Code Generation
@@
LLMs and EAs have both shown promise in automating code generation. LLMs can be trained on vast amounts of publicly available source code to gain a broad understanding of programming concepts and patterns [106, 107]. However, their generation abilities are limited by their training data distribution. EAs, on the other hand, are capable of open-ended search through program spaces. But traditional mutation operators used in GP struggle to propose high-quality changes in a way that mimics how human programmers intentionally modify code. The joining of EAs and LLMs has opened up more opportunities for code generation.
TABLE V: Summary of Applications Driven by Integrated Synergy of LLM and EA.
Category Subcategory Method Description Ref.
Code Generation Universal Code Generation ELM Universal method for code generation [26, 108]
WizardCoder Use Evol-Instruct to enhance the code generation [109]
Pinna et al. Improve LLM-generated code by Grammatical Evolution [110]
@@
Domain-specific Code Generation SEED Data cleaning tasks [50]
EUREKA Design reward in reinforcement learning [49]
EROM Design reward in reinforcement learning [111]
Zhang et al. Design reward in reinforcement learning [112]
Diffusion-ES Design reward in reinforcement learning [113]
GPT4AIGChip Design AI accelerator [51]
FunSearch For mathematical and algorithmic discovery [19]
L-AutoDA For decision-based adversarial attacks [114]
LLM-SR For scientific equation discovery from data [115]
Shojaee et al. Node importance scoring functions in complex networks [116]
Security in Code Generation DeceptPrompt Generates code containing specified vulnerabilities [117]
G3P with LLM Enhance code security [118]
Software Engineering Software Optimization Kang et al. Improves traditional genetic improvement [119]
Brownlee et al. Improves traditional genetic improvement [120]
Software Testing TitanFuzz Test case generation [18]
CodaMOSA Test case generation [121]
Software Project Planning SBSE Optimize selection of example sets [122]
Neural Architecture Search Representation Capability of LLM GPT-NAS Fine-tune GPT model to guide the NAS. [123]
Generation Capability of LLM LLMatic Generate code of architecture. [124]
Evoprompting Architecture generation through soft prompt tuning. [125]
Guided Evolution Use LLM to mutate and crossover the architecture code [126]
Others Use LLM as an operator to generate new architectures [127, 128]
GPTN-$S$ Employ GPT-4 to generate code of NAS algorithm [129]
Reasoning Capability of LLM Jawahar et al. Performance predictors of architecture. [130]
ReStruct Use LLM as predictor and selector of structure [131]
IV-A1 Universal Code Generation
Lehman et al. applied LLMs into GP, called evolution through large models (ELM) [26]. ELM utilizes a small number of initially hand-written programs as seeds, combining the MAP-Elites search algorithm [132] with an LLM-based intelligent mutation operation to generate a large amount of functionally rich programs as design examples. This method improves the search efficiency of GP algorithms since the LLM trained on code promote a more intelligent and effective mutation operator and reducing the probability of random mutations generating meaningless code. Considering the LLM may achieve different performance improvements in different domains, these programs are used as a training set to train LLMs in order to guide the LLM to output better code. Furthermore, by splicing new conditional inputs onto the LLM, it can teach the LLM to generate conditionally. This model will be fine-tuned through reinforcement learning to make the model able to generate appropriate outputs based on observed conditions. Bradley et al. proposed an open-source Python library based on ELM [108]. Luo et al. applied the Evol-Instruct introduced in Section III-A3 to generate a more complex and diverse training dataset [109] from the existing Code Alpaca instructions. They then fine-tuned the open-source code LLM StarCoder using this dataset, resulting in “WizardCoder” with the state-of-the-art performance on code generation. Different from the above studies, Pinna et al. did not directly apply evolutionary operators to code [110]. The code generated by LLMs are used as the initial individuals. And representing these individuals as abstract syntax trees, Grammatical Evolution algorithm continuously improve the code individuals and ultimately obtains the optimal code.
IV-A2 Domain-specific Code Generation
The combined approach of using LLM and EA has demonstrated practicality in various domains. These methods essentially leverage the iterative search framework of EA and the code generation and text understanding capabilities of LLM, with LLM acting as an evolutionary operator in the code evolution process. By iteratively improving the code, these methods enable the generation of algorithmic code by utilizing prompts that typically include contextual information such as environment source code, task descriptions, or example code. This approach has been applied to generate code for a wide range of tasks, including customized data cleaning methods [50], reward function design in reinforcement learning [49, 111, 112, 113], automated AI accelerator design [51], mathematical and scientific discoveries [19, 115], decision-based adversarial attacks [114], and criticality evaluation of nodes in complex networks [116].
In these research studies, additional innovations have further enhanced the collaborative effect of LLM and EA. For example, as introduced by Ma et al., the Evolution-driven Universal REward Kit for Agent (EUREKA) tracks the changing values of each component in the reward function throughout the reinforcement learning training process [49]. It monitors metrics like the value curve of an individual penalty term over time. EUREKA generates a list of these tracked changes in a reward feedback text provided to the language model for reference. This allows the method to guide the language model towards targeted improvements to the reward function design. Additionally, EUREKA also permits humans to provide purely textual descriptions of the strengths and weaknesses of the current reward function’s behavior as well as directions for desired behavioral changes. It takes these suggestions as context for the next round of search, generating reward functions that are better aligned with human preferences [49]. In Narin’s work, they additionally utilize the capabilities of a multimodal LLM, GPT-4V, to comprehend the visual information in the environment and assist in the generation process [111]. The FunSearch method designed by Romera et al. uses an island-based EA strategy to split the program database into multiple subpopulations that evolve in parallel [19]. This can effectively avoid local optima and maintain exploration diversity. Moreover, clustering programs based on their signatures can effectively identify programs with similar functionality but different code implementations, thus preserving more diverse programs. This balancing of exploitation and exploration enhances the algorithm’s ability to more comprehensively search the solution space and discover globally optimal solutions. The GPT for AI Generated Chip (GPT4AIGChip) proposed by Fu et al. introduces a demo-augmented prompt generator [51]. This generator can automatically select the two most relevant demonstrations from the demonstration library that correspond to the input design instruction. It augments the selected demonstrations into the prompt as contextual information to effectively guide the LLM towards generating more accurate code. The method proposed by Mao et al. optimized various details including population design, management, evaluation, and experimental design [116].
IV-A3 Security in Code Generation
@@
In addition to directly generating code, some studies have focused on the security of generated code. Wu et al. concerned about how to evaluate whether the generated code is vulnerable to attacks [117]. They proposed DeceptPrompt, which can generate adversarial natural language prefixes/suffixes to drive code generation models to produce functionally correct code containing specified vulnerabilities. DeceptPrompt continuously optimizes the prefixes and suffixes using GA, with LLM serving as the mutation operator. The fitness function is separately designed for the functionally correct part and the vulnerable part to guide the generated code to both retain functionality and contain the specified vulnerabilities. Empirical studies have shown that DeceptPrompt can successfully attack various popular code generation models. Tao et al. further considered methods to enhance the security of generated code [118]. They proposed a simple idea of combining LLM with Grammar-Guided GP (G3P) system [133], which enforces that any synthesized program adheres to Backus-Naur form (BNF) syntax, thus promoting the development/repair of incorrect programs and reducing the chances of security threats. Specifically, LLM is primarily employed to generate the initial code population in G3P, where the initial population is mapped to a predefined BNF syntax program before evolutionary search, allowing for improvements of these programs using program synthesis benchmarks. Afterwards, these programs will be handed over to G3P and evolved into better code.
IV-B Software Engineering
@@
Due to the promising performance of LLM and EA in code generation tasks, some studies have further applied them to practical problems in software engineering [134, 135], including subtakss software optimization, software testing, and software project planning.
IV-B1 Software Optimization
@@
Genetic Improvement (GI) [20] is a technique for automatically software optimization based on the ideas of evolutionary computation. Kang et al. leverage the advantages of LLM in code understanding and generation to improve the shortcomings of blind mutation in traditional GI [119]. The usage is similar to the code generation steps in Section IV-A, firstly letting the LLM improve the code for specific optimization objectives (such as time efficiency and memory consumption), then injecting the mutations generated by LLM into the candidate pool of GI, and continuously performing evolutionary operations until termination to obtain optimized code. Similarly, another study also enhanced the mutation operators of GI by leveraging LLMs to improve non-functional properties or fix functional bugs of software [120]. Brownlee et al. believed that the search space of GI is limited by the mutation operators it uses. By treating LLMs’ suggestions as additional mutation operators, they aimed to enrich the search space and thus obtain more successful mutation results (generated code).
IV-B2 Software Testing
@@
EAs have extensive applications in software testing techniques, where they are employed to search for effective test cases within the testing space to uncover errors and defects in software systems [136, 137]. Recently, some studies combine the code generation capability of LLMs and search capacity of EAs, and apply them to software testing. Deng et al. proposed TitanFuzz [18] to test for bugs in deep learning libraries. The core idea of TitanFuzz is to directly leverage LLMs that have been pretrained on tens of billions of code snippets, which implicitly learned the syntax and API constraints. It utilizes the generative ability of Codex [106] for seed generation and the infilling ability of InCoder [138] for mutation generation, to automatically produce a large number of input programs that meet the requirements for testing. The fitness function considers both the depth of the dataflow graph obtained from static analysis and the number of unique APIs called in the program, as well as the number of repeated API calls (as a penalty). This encourages the generation of programs with more complex API usage and richer API interactions. Lemieux et al. took a more direct approach [121], namely CodaMOSA. They first used conventional evolutionary search Many-Objective Sorting Algorithm (MOSA) [139] and monitored its coverage progress. When the algorithm reached a coverage plateau (state where further mutation of test cases finds it difficult to improve coverage [140, 141]), it would identify functions with low coverage in the module, and use these low-coverage functions as hints to request the LLM (e.g., CodeX [106]) to generate test cases. Afterwards, on the basis of the test cases generated by the LLM, the EA continued exploration. This method can significantly improve the coverage of test samples.
IV-B3 Software Project Planning
@@
Story points are a commonly used method in software project planning and cost estimation to gauge the amount of work required to complete a user story [142]. One way to leverage LLMs for this task is through few-shot learning, by providing some labeled examples of past user stories and their estimated points to improve the model’s ability to estimate new stories. However, the choice and number of examples can impact the estimation effectiveness. Tawosi et al. propose Search-Based Optimisation for Story Point Estimation (SBSE) to optimize the example selection through a multi-objective approach [122]. They use NSGA-II to simultaneously optimize three objective functions - the sum of absolute errors, the confidence interval of the error distribution, and the number of examples. This searches for the Pareto front of user story example sets. The GA explores the trade-off between estimation accuracy and complexity, as measured by the objective functions. The optimized Pareto front provides decision makers with different accuracy-complexity choice options when selecting an example set for few-shot learning with LLMs to estimate story points.
IV-C Neural Architecture Search (NAS)
LLMs and EAs also contribute to another field known as NAS [^4] Methods reviewed in this section differ from those presented in Section III-B. Approaches discussed in Section III-B primarily focus on LLM architecture search, and their techniques are based on EAs, whereas methods reviewed in this section leverage the synergistic combination of EAs and LLMs. Moreover, these NAS methods are more versatile and not limited to LLM architecture search alone, applicable to a broader range of NAS tasks.. Over the years, EAs have found widespread application in NAS [95, 96], revolutionizing the way we design and optimize neural networks. Their ability to simulate natural selection and iteratively improve architectures has proven invaluable in the quest for efficient and high-performing models. However, as we delve into the realm of cutting-edge advancements, a new player has emerged onto the NAS scene - LLMs. With their immense computational capacity and deep understanding of language and context, LLMs bring a fresh perspective to NAS. They possess the potential to offer novel insights and innovative solutions, empowering researchers and engineers to explore uncharted territories of network design. In the NAS applications based on LLM and EA, EAs are commonly employed to establish effective search frameworks, while LLMs leverage their unique abilities to contribute to NAS from diverse angles, including their representation capability, generation capability, and abundant prior knowledge.
IV-C1 With Representation Capability of Fine-tuned LLM
Yu et al. proposed a NAS method called GPT-NAS [123] that utilizes a GPT model [143] to guide the search of neural architectures. In GPT-NAS, neural architectures are encoded as inputs for the GPT model using a defined encoding strategy. Then the GPT model is pre-trained and fine-tuned on neural architecture datasets to introduce prior knowledge into the search process. In the architecture search, GPT-NAS uses an EA with operations like crossover and mutation to optimize individuals and obtain offspring with better performance. The fine-tuned GPT model is then used to effectively predict excellent new architectures based on previous structural information, thereby reconstructing the sampled architectures to guide and optimize the entire search process. In this method, the EA and GPT achieve complementary advantages. The EA excels in search and optimization, yet its direct application may face limitations due to constraints in variation capability and the vastness of the search space. This can pose challenges in effectively uncovering truly exceptional architectures. By reorganizing the structures on the basis of individuals evolved by GPT, richer prior knowledge can be introduced to optimize the network architectures, effectively reducing the actual search space and guiding the search towards high-performance areas.
IV-C2 With Code Generation Capability of General LLM
Nasir et al. leveraged the code generation ability of large models to help NAS [124]. The proposed LLMatic does not directly search in the structure representation space, but uses the code of neural network structures as search points. Specifically, LLMatic establishes two archives: a network archive that stores trained network models, and a prompt archive that stores prompts and temperature parameters for generating networks. LLMatic employs QD algorithms [144] to perform the search, utilizing the repeated process of generation, training and selection to continuously optimize the quality and diversity of networks. The large model CodeGen [145] is used to complete crossover and mutation in the evolutionary process: in the mutation operation, CodeGen makes subtle modifications to the selected network based on the mutation prompt. In the crossover operation, CodeGen partially matches the selected multiple networks based on the crossover prompt. Chen et al. took a similar approach by using the PALM [4] as the crossover and mutation operators for the NAS task in the evolutionary search process. Moreover, they further improved the LLM’s ability of generating candidate architectures through soft prompt tuning during NAS [125]. The Guided Evolution framework proposed by Morris et al. also uses LLM to mutate and crossover the code [126]. The main EAs adopted are SPEA-2 [146] and NSGA-II. SPEA-2 is used to retain elite individuals that performed well in the previous generation. NSGA-II is used to select individuals for crossover and mutation. At the same time, they proposed “Evolution of Thought” to continuously optimize the recommendations of LLM, which allows the LLM to propose improved modifications for new individuals based on the performance of past mutations, making the evolutionary process more data-driven and efficient. In addition to the above study, some papers [127, 128] do not explicitly mention using an EA framework, but employ LLM as an operator to generate new architectures. In addition to using LLM for architecture code generation, GPTN-SS employed GPT-4 to assist in generating code for a tensor network structure search algorithm [129]. Similar to other code generation methods discussed in Section IV-A, EA provides an iterative optimization framework, while LLM is utilized for evolutionary operations on the code.
IV-C3 With Reasoning Capability of LLM Based on Its Abundant Prior Knowledge
Unlike previous works that used LLM in the NAS search process, Jawahar et al. utilized LLM to build performance predictors [130]. Specifically, they designed predictor prompts (including task description, architecture parameter definitions, examples, etc.) to define the prediction task. The prompts and test architectures would be input into the LLM (such as GPT-4), and the LLM could make predictions based on the architecture knowledge it learned during pre-training by understanding the complex relationships between architectures and performances. Since LLMs learned architecture-related knowledge from a large amount of literature, and this type of predictor is designed simply without needing to train an over-parameterized network, the deployment cost is low. Chen et al. designed a novel reasoning meta-structure search framework called ReStruct to automatically discover meta-structures in heterogeneous information networks [131]. They developed a grammar translator that encodes meta-structures into natural language sentences, which facilitates LLMs’ understanding of structural patterns. Based on this, they designed a predictor and selector based on LLMs. The predictor adopts the idea of few-shot learning - it randomly samples some instances from the historical structure-performance records stored in previous searches, and uses these instances to prompt the LLM to predict the performance of new generated candidate structures. The selector will receive the prediction results of various candidate structures given by the predictor, and interact with the LLM to obtain selection suggestions. These modules are used in the selection process of GA, and combined with the designed insertion, grafting, and deletion operations to continuously optimize the meta-structure population and drive the search process. ReStruct also includes an explanation module based on chain-of-thought prompting, which can ultimately obtain semantically explainable and high-performance meta-structures.